# Reconciling aging and slip state evolutions from laboratory-derived canons of rate-and-state frictionF

### v2

@@ Appendix D Derivation of $B=B_{\rm heal}$ from other experimental requirements
However, to the authors’ surprise, the proposition $B=B_{\rm heal}$, in fact, can be derived as a logical consequence of some other experimental requirements (eqs. 26, 48, and 67), as long as we presume $\dot{\theta}=f(\theta,V)$ as we do when constructing the specific functional form of $f$. We show the proof below.
For simplicity, we consider $\theta\gg\theta_{\rm X}$, and associated approximations are denoted by $\simeq$. Note that neither $B$ nor $B_{\rm heal}$ matters for $\theta\ll\theta_{\rm X}$ because the strength $\Phi$, the observable, varies little with $\theta$, being approximately $\Phi_{\rm X}$. When $B\neq B_{\rm heal}$, $c_{*}$ is not an arbitrary constant, and then the convention $c_{*}=1$ (eq. 63) is not applicable. Hence, we keep $c_{*}$ as a variable in the subsequent derivation. The starting points are eq. (29), which follows the SS requirement of eq. (26), eq. (50), which is an identity, eqs. (55) and (56), which follow eq. (50) and the SHS requirement of eq. (48), and eq. (67), which is another SHS requirement.
@@
$c_{*}^{-1}f_{\rm SHS|NSC}\simeq(B_{\rm heal}/B)(c_{*}^{-1}\theta_{\rm SHS})^{1-B/B_{\rm heal}}.$ (D.98)
@@
$f_{\rm SHS|NSC}=\frac{B_{\rm heal}}{B}\frac{L}{L_{\rm heal}}\left(\frac{V_{\rm prior}\theta_{\rm SHS}}{L}\right)^{1-B/B_{\rm heal}}.$ (D.99)
@@
Thus, even if we do not assume $B=B_{\rm heal}$ as an experimental requirement (eq. 49), the development of the evolution law is not affected, as long as it presumes an evolution law depending only on the rate $V$ and state $\theta$ with temporal locality, $\dot{\theta}=f(\theta,V)$. Besides, we note that the above analysis does not exclude the possibility that temporally local evolution laws with $B\neq B_{\rm heal}$ become possible if the evolution laws include responses to factors other than $V$ and $\theta$ (e.g., normal and shear stresses; Linker & Dieterich, 1992; Nagata et al., 2012).
@@
 [^1] Proofreading memorandum: In the section titles, the latex ref command seems not working in GJI formats. If 93 and 95 in the title of Appendix E are different from equation numbers of (93) and (95), please replace them in the title manually.
@@
Therefore, what we aim to prove here is the proposition that $\psi_{\rm cBase}$ decaying as strongly as eq. (100) tells already satisfies eq. (83) except for the range $|\Omega-1|\lesssim|\beta\ln\beta|$. Since eq. (E.6) is the same as eq. (100) for $\Omega\ll\beta$, we only need to prove it for $\Omega\gg\beta$, where the concerned proposition (if eq. 100 is true then eq. E.6 is true) is reduced to the following: if $\psi_{\rm cBase}=\mathcal{O}[(L/D_{\rm c})\beta\ln\beta]$, then $\psi_{\rm cBase}=o[(L/D_{\rm c})\Omega\ln\Omega]$ except for $|\Omega-1|\lesssim|\beta\ln\beta|$. Namely,
@@
We begin with the range $\beta\ll\Omega<1/e$. Note $\beta\ll\Omega<\sqrt{\beta}<1/e$ in this appendix because $\sqrt{\beta}\ll 1$ and $\beta<1/e^{2}$ are assumed. Then we can consider an $\Omega$ range $\beta\ll\Omega<\sqrt{\beta}$ within $\Omega<1/e$. For $\beta\ll\Omega<\sqrt{\beta}$, $\Omega/\beta<1/\sqrt{\beta}$ holds. Then, taking its logarithm, we see $\ln(\Omega/\beta)<(1/2)|\ln\beta|$, which is followed by
@@
$\frac{\beta\ln\beta}{\Omega\ln\Omega}<2\beta/\Omega=\mathcal{O}(\beta/\Omega)=o(1).$ (E.11)
@@
Equations (E.13), (E.15), and (E.17) combined are equivalent to eq. (E.7), and hence eq. (E.6) has been proven for $\psi_{\rm cBase}$ that satisfies eq. (100).
@@
In this subsection §E.2, we consider a special case where the decreasing function $\psi_{\rm cBase}$ additionally satisfies all the conditions of eqs. (E.3)–(E.5). Equation (E.3) constrains the decreasing function $\psi_{\rm cBase}$ to decay sufficiently below $(L/D_{\rm c})\beta$ by $\Omega=1/2$. This stipulation does not necessarily have to be made at $\Omega=$$1/2$”. In fact, the following proof of eq. (E.2) works even when eq. (E.3) is replaced by the same constraint at $\Omega=1-|\beta^{\prime}\ln\beta^{\prime}|$, where $\beta^{\prime}$ is a constant such that $\beta\ll\beta^{\prime}\ll 1$:
$\psi_{\rm cBase}(1-|\beta^{\prime}\ln\beta^{\prime}|)\ll(L/D_{\rm c})\beta,$ (E.18)
because the role of eq. (E.3) in the proof is to guarantee that $\psi_{\rm cBase}$ is sufficiently small at the minimum $\Omega$ in $\beta\lesssim|\Omega-1|\lesssim|\beta\ln\beta|$. In the proof below, we adopt eq. (E.3) just for brevity.
Equation (E.2) represents the tolerance of $\epsilon_{\Omega}^{*}$ if eq. (83) holds outside the $\epsilon_{\Omega}^{*}$-neighborhood of $\Omega=1$ with $\epsilon_{\Omega}^{*}$ satisfying eq. (E.2). Thus, we below show that eq. (83 holds for $|\Omega-1|\gg(D_{\rm c}/L)\psi_{\rm cBase}(1)$ given that $\psi_{\rm cBase}$ satisfies the additional assumptions eqs. (E.3)–(E.5), atop eq. (100).
@@
We evaluate eq. (83) separately for the range $\beta\lesssim|\Omega-1|\lesssim|\beta\ln\beta|$ and the other range $(D_{\rm c}/L)\psi_{\rm cBase}(1)\ll|\Omega-1|\ll\beta$ (if exists). We begin with the former. The assumption of eq.  (E.3), together with the monotonic decrease of $\psi_{\rm cBase}$, guarantees the following inequality for $\Omega\geq 1/2$, which includes the currently examined range $\beta\lesssim|\Omega-1|\lesssim|\beta\ln\beta|$:
@@
Next, we work on the other range $(D_{\rm c}/L)\psi_{\rm cBase}(1)\ll|\Omega-1|\ll\beta$. When $|\Omega-1|\ll\beta$, we have
$|\Omega\ln\Omega|=|\Omega-1|+\mathcal{O}[(\Omega-1)^{2}].$ (E.22)
Meanwhile, eqs. (E.4) and (E.5) guarantee
@@
From eqs. (E.22) and (E.23), for $\Omega$ such that $(D_{\rm c}/L)\psi_{\rm cBase}(1)\ll|\Omega-1|\ll\beta$, if exists, we have
$|\Omega\ln\Omega|=|\Omega-1|+\mathcal{O}[(\Omega-1)^{2}]\gg(D_{\rm c}/L)\psi(1)=(D_{\rm c}/L)\psi(\Omega)[1+\mathcal{O}(|\Omega-1|/\beta)],$ (E.24)
@@
Besides, we note that even without the assumption of eqs. (E.4) and (E.5), the assumption of eq. (E.3) is enough to relax the lower bound of $\epsilon_{\Omega}^{*}$ to $\beta$ from $|\beta\ln\beta|$ in eq. (E.1), although $\beta$ and $|\beta\ln\beta|$ are of almost the same order unless $\beta$ is very small.
@@
Before using eq. (E.2), we need to confirm that it is applicable to this case with $\psi_{\rm cBase}=\exp(-\Omega/\beta)$. From $\beta^{-1}\exp(-(0.5\beta)/\beta)\ll 1$ and $d\exp(-x)/dx=\mathcal{O}[\exp(-x)]$, it is easy to see that $\psi_{\rm cBase}=\exp(-\Omega/\beta)$ satisfies eqs. (E.3)–(E.5), the additional requirements on $\psi_{\rm cBase}$ for eq. (E.2) to be applicable. The remaining job is to evaluate if eq. (90) (eq. 100 for $L\sim D_{\rm c}$) holds for $\psi_{\rm cBase}=\exp(-\Omega/\beta)$. For this purpose, we consider a more general case where $\psi_{\rm cBase}=\exp[-(\Omega/\beta)^{n}]$ and determine the range of $n(>0)$ satisfying eq. (90); or equivalently, we determine the range of $n>0$ such that
@@
As the present paper takes a position that $\beta\lesssim 0.01$ has been constrained by existing experiments (eq. 85), $\exp(-1/\beta)$ is very small, being $10^{-43.24...}$ for $\beta=0.01$. Hence, we can use very small $\epsilon^{*}_{\Omega}$ so that the $\epsilon^{*}_{\Omega}$-neighborhood of the steady state, namely, the range $|\Omega-1|<\epsilon^{*}_{\Omega}$ in which $\psi_{\rm c}$ has to satisfy eq. (91), can be taken so narrow that this range becomes numerically null, which is good news. Nevertheless, it is still necessary to set $\epsilon_{\Omega}^{*}$ appropriately for given $\beta$, according to eq. (E.30). A convenient setting is $\epsilon_{\Omega}^{*}=\sqrt{e^{-1/\beta}}$, which necessarily satisfies eq. (E.30) as far as $\beta\ll 1$.
@@ Appendix F Construction of $\beta$ for a given $\psi_{\rm c}$
In §4.2, we have derived eq. (83) as a constraint that $\psi_{\rm c}(\Omega)$ ($=f_{t}$) must conform. When compared with eq. (84), eq. (83) can also be read as the expression of the scale of $\Omega$ around which the aging-slip switch must occur. When we constructed $\psi_{\rm c}$ for a given $\beta$ in §4.3, we considered the functions of $\Omega$ that have $\beta$ as the scale parameter (more specifically, unary functions of $\Omega/\beta$). Alternatively, eq. (83) allows the opposite approach where we adjust the value of $\beta$ so that the canons can be reproduced with a given $\psi_{\rm c}(\Omega)$. Obviously, such an approach is necessary when we consider $\psi_{\rm c}$ forms that do not involve $\beta$ as an explicit parameter, as in the case of eq. (103) discussed in §4.4. Also, the inversion of $\beta$ from experimentally obtained state evolution (or from its time derivative, $f=f_{t}+f_{\delta}V$) belongs to the latter approach. In the below, we present a general procedure to construct $\beta$ for a given $\psi_{\rm c}$, followed by application to two examples of $\psi_{\rm c}(\Omega)$ discussed in §4.4, where this procedure is necessary.
As discussed above, the value of $\beta$ in eq. (83) is the scale of $\Omega$ around which the aging-slip switch occurs (eq. 84). Since eq. (84) does not specify the behavior for $\Omega\sim\beta$, the steepness of the aging-slip crossover is unknown. Then, some 1-order ambiguity is intended in the concept of $\beta$. Nevertheless, from the viewpoint of eventually writing down a concrete evolution law, it is desirable to assign a single value to $\beta$. In this sense, we propose, as a practical procedure, to define a representative value $\beta^{\prime}$ to be the value of $\Omega$ such that
@@
As mentioned above, somewhat different values also function as $\beta$ due to the unconstrained transient behavior of the aging-slip switch, but $\beta^{\prime}$ obtained by the above procedure necessarily satisfies eq. (83) whatever the crossover is, implying that the proposed procedure is robust. Actually, the value of $\beta$ appears as $\ln\beta$ in $\Phi/B$ so that this 1-order ambiguity in the $\beta$ value is just within an $\mathcal{O}(B)$ factor in the strength evolution law.
@@
Since this $\psi_{\rm cBase}$ does not have the parameter $\beta$, this is an obvious case where we have to determine $\beta$ for the given $\psi_{\rm c}$. The $f_{\delta}V$ term in this example is already fixed by eq. (72). Then, eq. (F.2) becomes
@@
The said $\psi_{\rm cBase}$ of eq. (F.3) satisfies eq. (F.5), so we can use eq. (E.1) to obtain the associated tolerance of $\epsilon_{\Omega}^{*}$. Using the $\beta=\beta^{\prime}$ value of eq. (F.4), eq. (E.1) reduces to

### v3

@@ Appendix D Derivation of $B=B_{\rm heal}$ from other empirical formulae
However, to the authors’ surprise, the proposition $B=B_{\rm heal}$, in fact, can be derived as a logical consequence of some other empirical formulae (eqs. 26, 48, and 67), as long as we presume $\dot{\theta}=f(\theta,V)$ as we do when constructing the specific functional form of $f$. We show the proof below.
For simplicity, we consider $\theta\gg\theta_{\rm X}$, and associated approximations are denoted by $\simeq$. Note that neither $B$ nor $B_{\rm heal}$ matters for $\theta\ll\theta_{\rm X}$ because the strength $\Phi$, the observable, varies little with $\theta$, being approximately $\Phi_{\rm X}$. When $B\neq B_{\rm heal}$, $c_{*}$ is not an arbitrary constant, and then the convention $c_{*}=1$ (eq. 63) is not applicable. Hence, we keep $c_{*}$ as a variable in the subsequent derivation. The starting points are eq. (29), which follows the SS requirement of eq. (26), eq. (50), which is an identity, eqs. (55) and (56), which follow eq. (50) and the SHS requirement of eq. (48), and eq. (67), which is another empirical formula of SHS tests.
@@
$c_{*}^{-1}f_{\rm SHS|NSC}\simeq(B_{\rm heal}/B)(c_{*}^{-1}\theta_{\rm SHS|NSC})^{1-B/B_{\rm heal}}.$ (D.98)
@@
$f_{\rm SHS|NSC}=\frac{B_{\rm heal}}{B}\frac{L}{L_{\rm heal}}\left(\frac{V_{\rm prior}\theta_{\rm SHS|NSC}}{L}\right)^{1-B/B_{\rm heal}}.$ (D.99)
@@
Thus, even if we do not set forth $B=B_{\rm heal}$ as an experimental requirement (eq. 49), the development of the evolution law from the canons and $t_{\rm c}=L/V_{\rm prior}$ (eq. 67) is not affected, as long as it presumes an evolution law depending only on the rate $V$ and state $\theta$ with temporal locality, $\dot{\theta}=f(\theta,V)$. However, we note that eq. (67) is yet to be scrutinized under the NSC, and the above analysis does not exclude the possibility that temporally local evolution laws with $B\neq B_{\rm heal}$ become possible if the evolution laws include responses to factors other than $V$ and $\theta$ (e.g., normal and shear stresses; Linker & Dieterich, 1992; Nagata et al., 2012).
@@
Therefore, what we aim to prove here is the proposition that $\psi_{\rm cBase}$ decaying as strongly as eq. (100) tells already satisfies eq. (83) except for the range $|\Omega-1|\lesssim|\beta\ln\beta|$. Since eq. (E.6) is the same as eq. (100) for $\Omega\ll\beta$, we only need to prove it for $\Omega\gg\beta$, where the concerned proposition (if eq. 100 is true then eq. E.6 is true) is reduced to the following: if $\psi_{\rm cBase}=\mathcal{O}[(L/D_{\rm c})\beta\ln\beta]$ (eq. 100), then $\psi_{\rm cBase}=o[(L/D_{\rm c})\Omega\ln\Omega]$ except for $|\Omega-1|\lesssim|\beta\ln\beta|$ (eq. E.6). Namely,
@@
We begin with the range $\beta\ll\Omega<1/e$. Note, in this appendix E, $\beta\ll\sqrt{\beta}<1/e$ holds because $\sqrt{\beta}\ll 1$ and $\beta<1/e^{2}$ are assumed. Then we can consider an $\Omega$ range $\beta\ll\Omega<\sqrt{\beta}$ within $\Omega<1/e$. For $\beta\ll\Omega<\sqrt{\beta}$, $\Omega/\beta<1/\sqrt{\beta}$ holds. Then, taking its logarithm, we see $\ln(\Omega/\beta)<(1/2)|\ln\beta|$, which is followed by
@@
$\frac{\beta\ln\beta}{\Omega\ln\Omega}<2\beta/\Omega=o(1).$ (E.11)
@@
Equations (E.13), (E.15), and (E.17) combined are equivalent to eq. (E.7), and hence eq. (E.6) has been proven for $\psi_{\rm cBase}$ that satisfies eq. (100). [Note: Strictly speaking, what the calculus leading to eqs. (E.12) and (E.15) shows is $|\Omega\ln\Omega|\gg|\beta\ln\beta|$ for $\Omega\gg 2\beta$, or $|\Omega\ln\Omega|\gg|\beta\ln\beta|/2$ for $\Omega\gg\beta$, not being $|\Omega\ln\Omega|\gg|\beta\ln\beta|$ for $\Omega\gg\beta$. However, this is not a problem for the present purpose because we do not distinguish $\Omega\sim\beta$ and $\Omega\sim 2\beta$ in constructing the evolution law from eq. (100).]
@@
In this subsection §E.2, we consider a special case where the decreasing function $\psi_{\rm cBase}$ additionally satisfies all the conditions of eqs. (E.3)–(E.5). Equation (E.3) constrains the decreasing function $\psi_{\rm cBase}$ to decay sufficiently below $(L/D_{\rm c})\beta$ by $\Omega=1/2$. This stipulation does not necessarily have to be made at $\Omega=$$1/2$”. In fact, the following proof of eq. (E.2) works even when eq. (E.3) is replaced by the same constraint at $\Omega=1-|\tilde{\beta}\ln\tilde{\beta}|$, where $\tilde{\beta}$ is a constant such that $\beta\ll\tilde{\beta}\ll 1$:
$\psi_{\rm cBase}(1-|\tilde{\beta}\ln\tilde{\beta}|)\ll(L/D_{\rm c})\beta,$ (E.18)
because the role of eq. (E.3) in the proof is to guarantee that $\psi_{\rm cBase}$ is sufficiently small at the minimum $\Omega$ in $\beta\lesssim|\Omega-1|\lesssim|\beta\ln\beta|$ (in eq. E.19). In the proof below, we adopt eq. (E.3) just for brevity.
Proving eq. (E.2) as (a subset of) the tolerance of $\epsilon_{\Omega}^{*}$ is equivalent to showing that eq. (83) holds outside the $\epsilon_{\Omega}^{*}$-neighborhood of $\Omega=1$ with $\epsilon_{\Omega}^{*}$ satisfying eq. (E.2). Thus, we below show that eq. (83 holds for $|\Omega-1|\gg(D_{\rm c}/L)\psi_{\rm cBase}(1)$ given that $\psi_{\rm cBase}$ satisfies the additional assumptions eqs. (E.3)–(E.5), atop eq. (100).
@@
We evaluate eq. (83) firstly for the range $\beta\lesssim|\Omega-1|\lesssim|\beta\ln\beta|$ and secondly for the remaining range $(D_{\rm c}/L)\psi_{\rm cBase}(1)\ll|\Omega-1|\ll\beta$ (if exists). The assumption of eq.  (E.3), together with the monotonic decrease of $\psi_{\rm cBase}$, guarantees the following inequality for $\Omega\geq 1/2$, which includes the currently examined range $\beta\lesssim|\Omega-1|\lesssim|\beta\ln\beta|$:
@@
Next, we work on the other range $(D_{\rm c}/L)\psi_{\rm cBase}(1)\ll|\Omega-1|\ll\beta$. Equations (E.4) and (E.5) guarantee
@@
From eq. (E.22), for $\Omega$ such that $(D_{\rm c}/L)\psi_{\rm cBase}(1)\ll|\Omega-1|\ll\beta$, if exists, we have
$|\Omega\ln\Omega|=|\Omega-1|+\mathcal{O}[(\Omega-1)^{2}]\gg(D_{\rm c}/L)\psi_{\rm cBase}(1)=(D_{\rm c}/L)\psi_{\rm cBase}(\Omega)[1+\mathcal{O}(|\Omega-1|/\beta)],$ (E.23)
@@
Besides, we note that even without the assumption of eqs. (E.4) and (E.5), the assumption of eq. (E.3) is enough to relax the lower bound of $\epsilon_{\Omega}^{*}$ to $\beta$ from $|\beta\ln\beta|$ in eq. (E.1) (via eq. E.21), although $\beta$ and $|\beta\ln\beta|$ are of almost the same order unless $\beta$ is very small.
@@
Before using eq. (E.2), we need to confirm that it is applicable to this case with $\psi_{\rm cBase}=\exp(-\Omega/\beta)$. From $\exp(-0.5/\beta)\ll\beta$ and $d\exp(-x)/dx=\mathcal{O}[\exp(-x)]$, it is easy to see that $\psi_{\rm cBase}=\exp(-\Omega/\beta)$ satisfies eqs. (E.3)–(E.5), the additional requirements on $\psi_{\rm cBase}$ for eq. (E.2) to be applicable. The remaining job is to evaluate if eq. (90) (eq. 100 for $L\sim D_{\rm c}$) holds for $\psi_{\rm cBase}=\exp(-\Omega/\beta)$. For this purpose, we consider a more general case where $\psi_{\rm cBase}=\exp[-(\Omega/\beta)^{n}]$ and determine the range of $n(>0)$ satisfying eq. (90). Or equivalently, we determine the range of $n>0$ such that
@@
Since $\beta\lesssim 0.01$ has been constrained by existing experiments (eq. 85), $\exp(-1/\beta)$ is very small, being $10^{-43.24...}$ for $\beta=0.01$. Hence, we can use very small $\epsilon^{*}_{\Omega}$ so that the $\epsilon^{*}_{\Omega}$-neighborhood of the steady state, namely, the range $|\Omega-1|<\epsilon^{*}_{\Omega}$ in which $\psi_{\rm c}$ has to satisfy eq. (91), can be taken so narrow that this range becomes numerically null, which is good news. Nevertheless, it is still necessary to set $\epsilon_{\Omega}^{*}$ appropriately for given $\beta$, according to eq. (E.29). A convenient setting is $\epsilon_{\Omega}^{*}=\sqrt{e^{-1/\beta}}$, which necessarily satisfies eq. (E.29) as far as $\beta\ll 1$.
@@ Appendix F Construction of $\beta$ for a given $\psi_{\rm c}$
In §4.2, we have derived eq. (83) as a constraint that $\psi_{\rm c}(\Omega)$ ($=f_{t}$) must conform. When compared with eq. (84), eq. (83) can also be read as the expression of the scale of $\Omega$ around which the aging-slip switch must occur. When we constructed $\psi_{\rm c}$ for a given $\beta$ in §4.3, we implemented $\psi_{\rm c}$ for give $\beta$, which worked as the scale parameter in a unary function of $\Omega/\beta$. Alternatively, eq. (83) allows the opposite approach, introducing $\beta$ for given $\psi_{\rm c}$, where we adjust the value of $\beta$ so that the canons can be reproduced with a given $\psi_{\rm c}(\Omega)$. Obviously, such an approach is necessary when we consider $\psi_{\rm c}$ forms that do not involve $\beta$ as an explicit parameter, as in the case of eq. (103) discussed in §4.4. Also, the inversion of $\beta$ from experimentally obtained state evolution (or from its time derivative, $f=f_{t}+f_{\delta}V$) belongs to the latter approach. In the below, we present a general procedure to construct $\beta$ for a given $\psi_{\rm c}$, followed by application to examples of $\psi_{\rm c}(\Omega)$ discussed in §4.4, where this procedure is necessary.
As discussed above, the value of $\beta$ in eq. (83) is the scale of $\Omega$ around which the aging-slip switch occurs (eq. 84). Since eq. (84) does not specify the behavior for $\Omega\sim\beta$, the steepness of the aging-slip switch—namely, whether the switch is a crossover or a transition—is unknown. Then, some 1-order ambiguity is intended in the concept of $\beta$. Nevertheless, from the viewpoint of eventually writing down a concrete evolution law, it is desirable to assign a single value to $\beta$. In this sense, we propose, as a practical procedure, to define a representative value $\beta^{\prime}$ to be the value of $\Omega$ such that
@@
As mentioned above, somewhat different values also function as $\beta$ due to the unconstrained transient behavior of the aging-slip switch, but $\beta^{\prime}$ obtained by the above procedure necessarily satisfies eq. (83) whatever the switch is, implying that the proposed procedure is robust. Actually, the value of $\beta$ appears as $\ln\beta$ in $\Phi/B$ so that this 1-order ambiguity in the $\beta$ value is just within an $\mathcal{O}(B)$ factor in the strength evolution law.
@@
Since this $\psi_{\rm cBase}$ does not involve the parameter $\beta$, this is an obvious case where we have to determine $\beta$ for the given $\psi_{\rm c}$. The $f_{\delta}V$ term in this example is already fixed by eq. (72). Then, eq. (F.2) becomes
@@
The said $\psi_{\rm cBase}$ of eq. (F.3) satisfies eq. (F.5) (i.e., eq. 100 for eq. F.4), so we can use eq. (E.1) to obtain the associated tolerance of $\epsilon_{\Omega}^{*}$. Using the $\beta=\beta^{\prime}$ value of eq. (F.4), eq. (E.1) reduces to